### v1

# Breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation

Abstract
In this paper, we systematically investigate the intricate dynamics of the breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov-Ivanov equation. We discuss the transition conditions of the breather-to-soliton and obtain different types of nonlinear converted waves, including the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution. Meanwhile, the interactions among the above nonlinear converted waves are explored by choosing appropriate parameters. Furthermore, we derive the double-pole breather-to-soliton transitions and apply the asymptotic analysis method to analyze the dynamics of the asymptotic solitons for the double-pole anti-dark soliton.
keywords:
Higher-order modified Gerdjikov–Ivanov equation, Breather-to-soliton transitions, Nonlinear wave interactions, Double-pole converted waves, Asymptotic analysis
[…]
[1]organization=School of Mathematical Science, addressline=Beihang University, city=Beijing, postcode=102206, country=China
1 Introduction
It is well-known that solitons, breathers, and rogue waves represent fundamental nonlinear wave phenomena, each exhibiting distinct dynamic behaviors across diverse physical domains. They have attracted extensive attention due to their emergence in numerous nonlinear models involving various fields such as plasmas physics, nonlinear optics, hydrodynamics and Bose-Einstein condensations[2, 3, 4, 5, 6].
Breathers, which describe the growth of the perturbations in finite continuous backgrounds, contain two different types, namely Kuznetsov–Ma breathers and Akhmediev breathers [7, 8]. Recently, the state transition mechanisms between the breathers (or rogue waves) and solitons have been explored in various nonlinear models. Abundant converted waves including multi-peak solitons, periodic waves and anti-dark solitons have been presented starting from the breathers or rogue waves by an appropriate transitions condition. The first case involves the nonlinear models featuring high-order terms that contribute to the state transitions. The mixed nonlinear Schrödinger equation (NLS)[9], fifth-order NLS[10], higher-order generalized NLS[11] and six-order NLS equation[12] have been successively discovered to possess various converted waves and nonlinear wave interactions. The Hirota equation[13, 14, 15], as an extended NLS-type equation with high-order dispersion terms, also exhibits the breather-to-soliton transition phenomenon. Meanwhile, Sasa–Satsuma equation[16] and the high-order Chen–Lee–Liu equation have also been considered[17]. Another case occurs in coupled equations describing multi-field interactions which facilitate the state transitions. The transitions of the breather solution have been progressively discussed in the coupled Hirota equation[18, 19], multi-component AB system[20, 21, 22], coupled Fokas–Lenells system[23] and coupled extended modified Korteweg-de Vries system[24].
In this paper, we focus on the the higher-order modified Gerdjikov–Ivanov equation (HMGI),
$iq_{t}$ $+2iq_{xxx}+\left[(16\delta+4)|q|^{2}-4\delta\right]q_{xx}+(4\delta+1)q^{2}q_{xx}^{*}+(12\delta+3)q^{*2}q_{x}^{2}$
$+\left[(20\delta+2)qq_{x}^{*}-\left(48\delta^{2}+30\delta+\frac{15}{4}\right)i|q|^{4}\right.\left.+\left(24\delta^{2}+6\delta\right)i|q|^{2}-2i\delta^{2}\right]q_{x}$
$-\left[(24\delta^{2}+6\delta)|q|^{2}q^{2}-8\delta^{2}q^{2}\right]iq_{x}^{*}-\left(16\delta^{3}+24\delta^{2}+\frac{15}{2}\delta+\frac{5}{8}\right)|q|^{6}q$
$+\left(16\delta^{3}+12\delta^{2}+\frac{3}{2}\delta\right)|q|^{4}q-\left(4\delta^{3}+\delta^{2}\right)|q|^{2}q+2\delta^{3}q=0.$ (1)
where $\delta$ is an arbitrary real number which imposes a significant influence on the self-steepening effect in transmission which is the self-frequency shift of Raman scattering. Guan, Yang and Meng et.al. obtained the high-order rogue wave solutions and their dynamic properties for the HMGI equation by generalized Darboux transformation method[25]. The soliton, breather and new rogue wave solutions of the HMGI euation are reached by Zhao, Wang and Yang through Darboux transformation method[26]. The multi-positon solutions, interaction solutions and the modulation instability for the HMGI equation are exihited by Zhao, Wang and Zhang[27].
To our knowledge, the state transitions of breather-to-soliton for the HMGI equation have not been explored yet. This paper is organized as follows: in section 2, we provide the state transition conditions by analyzing the explicit expression of the breather solution and obtain different nonlinear converted waves including W-shaped soliton, M-shape soliton, anti-dark soliton, multi-peak soliton and periodic waves. In section 3, the interactions between the above various nonlinear converted waves are investigated. In section 4, double-pole breather-to-soliton transition solutions are derived from the double-pole breather solutions of the HMGI equation. In addition, the asymptotic analysis is given for the double-pole anti-dark soliton solution to obtain the explicit expression of the asymptotic solitons.
2 Breather-to-soliton transitions
In this section, we discuss the existence conditions for several types of nonlinear converted waves by analyzing the explicit expression of the first-order solution on the plane wave background. The Lax pair for Eq.(1) have given in Ref.[27] as follows,
$\Phi_{x}=P\Phi,~{}\Phi_{t}=Q\Phi,$ (2)
where $\Phi=(\psi,\phi)^{\mathrm{T}}$ is the complex vector function and
$P=\begin{pmatrix}\frac{1}{2}\mathrm{i}\lambda^{2}+\mathrm{i}\delta|q|^{2}&\frac{1}{2}\lambda q\\ -\frac{1}{2}\lambda q^{*}&-\frac{1}{2}\mathrm{i}\lambda^{2}-\mathrm{i}\delta|q|^{2}\end{pmatrix},~{}~{}Q=\begin{pmatrix}Q_{1}&Q_{2}\\ Q_{3}&-Q_{1}\end{pmatrix},$
with
$Q_{1}$ $=\frac{1}{2}iq^{*}q_{xx}+\frac{1}{2}iqq_{xx}^{*}+\frac{1}{16}\left[-8iq_{x}^{*}+(488+12)|q|^{2}q-\left(8\lambda^{2}+16\delta\right)q^{*}\right]q_{x}$
$-3\left[q^{*}\left(\delta+\frac{1}{4}\right)q-\frac{\lambda^{2}}{6}-\frac{\delta}{3}\right]qq_{x}^{*}-6i|q|^{6}\left(\delta^{2}+\frac{\delta}{2}+\frac{5}{96}\right)$
$+4i\left(\delta+\frac{3}{16}\right)\left(\frac{\lambda^{2}}{2}+\delta\right)|q|^{4}-\frac{i\left(\lambda^{2}+\delta\right)^{2}|q|^{2}}{2}+i\lambda^{6}+2i\lambda^{4}\delta+i\lambda^{2}\delta^{2}+i\delta^{3},$
$Q_{2}$ $=2\lambda\bigg{[}-\frac{q_{xx}}{2}+\bigg{[}3i\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}i\lambda^{2}-i\delta\bigg{]}q_{x}+\bigg{[}i\delta qq_{x}^{*}+\frac{(\lambda^{2}+\delta)^{2}}{2}$
$\quad-2\bigg{(}\delta+\frac{1}{4}\bigg{)}\bigg{(}\frac{\lambda^{2}}{2}+\delta\bigg{)}|q|^{2}+2\bigg{(}\delta^{2}+\frac{3}{4}\delta+\frac{3}{32}\bigg{)}|q|^{4}\bigg{]}q\bigg{]},$
$Q_{3}$ $=2\lambda\bigg{[}\frac{q_{xx}^{*}}{2}+\bigg{[}3i\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}i\lambda^{2}-i\delta\bigg{]}q_{x}^{*}+\bigg{[}i\delta^{*}qq_{x}-\frac{(\lambda^{2}+\delta)^{2}}{2}$
$\quad+2\bigg{(}\delta+\frac{1}{4}\bigg{)}\bigg{(}\frac{\lambda^{2}}{2}+\delta\bigg{)}|q|^{2}-2\bigg{(}\delta^{2}+\frac{3}{4}\delta+\frac{3}{32}\bigg{)}|q|^{4}\bigg{]}q^{*}\bigg{]},$
where $\lambda$ is the complex spectral parameter and the zero curvature equation $P_{t}-Q_{x}+PQ-QP=0$ generate original Eq.(1).
We introduce the zeed solution $q[0]=ce^{\mathrm{i}(ax+bt)}$ to Eq.(2) with
$b=$ $-16\delta^{3}c^{6}-24\delta^{2}c^{6}-\frac{15}{2}c^{6}\delta-\frac{5}{8}c^{6}+16\delta^{3}c^{4}+24\delta^{2}c^{4}a+12\delta^{2}c^{4}+24c^{4}\delta a$
$+\frac{3}{2}c^{4}\delta+\frac{15}{4}c^{4}a-4\delta^{3}c^{2}-16\delta^{2}c^{2}a-\delta^{2}c^{2}-12c^{2}a^{2}\delta-6c^{2}\delta a-6c^{2}a^{2}$
$+2a^{3}+4a^{2}\delta+2\delta^{2}a+2\delta^{3}.$
According to the Darboux transformation given in [26], we assume $\Phi_{j}=(\psi_{j},\phi_{j})^{\mathrm{T}},~{}j=1,\cdots,n$ are the eigenfunctions under $\lambda=\lambda_{j},~{}j=1,\cdots,n$, then we can obtain the $n$-order solution as follows,
$q[n]=q[0]-\dfrac{2\mathrm{i}|\Omega_{1}|}{|\Omega_{2}|},$ (3)
with
$\Omega_{2}=\begin{pmatrix}\lambda_{1}^{2n-2}\psi_{1}&\lambda_{1}^{2n-4}\psi_{1}&\cdots&\psi_{1}&\lambda_{1}^{2n-1}\phi_{1}&\lambda_{1}^{2n-3}\phi_{1}&\cdots&\lambda_{1}\phi_{1}\\ \lambda_{2}^{2n-2}\psi_{2}&\lambda_{2}^{2n-4}\psi_{2}&\cdots&\psi_{2}&\lambda_{2}^{2n-1}\phi_{2}&\lambda_{2}^{2n-3}\phi_{2}&\cdots&\lambda_{2}\phi_{2}\\ \vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ \lambda_{n}^{2n-2}\psi_{n}&\lambda_{n}^{2n-4}\psi_{n}&\cdots&\psi_{n}&\lambda_{n}^{2n-1}\phi_{n}&\lambda_{n}^{2n-3}\phi_{n}&\cdots&\lambda_{n}\phi_{n}\\ \lambda_{1}^{*2n-2}\phi_{1}^{*}&\lambda_{1}^{*2n-4}\phi_{1}^{*}&\cdots&\phi_{1}^{*}&-\lambda_{1}^{*2n-1}\psi_{1}^{*}&-\lambda_{1}^{*2n-3}\psi_{1}^{*}&\cdots&-\lambda_{1}^{*}\psi_{1}^{*}\\ \vdots&\vdots&\cdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ \lambda_{n}^{*2n-2}\phi_{n}^{*}&\lambda_{n}^{*2n-4}\phi_{n}^{*}&\cdots&\phi_{n}^{*}&-\lambda_{n}^{*2n-1}\psi_{n}^{*}&-\lambda_{n}^{*2n-3}\psi_{n}^{*}&\cdots&-\lambda_{n}^{*}\psi_{n}^{*}\end{pmatrix},$
and $\Omega_{1}$ is given by replacing (n+1)-th column in $\Omega_{2}$ with
$(-\lambda_{1}^{2n}\psi_{1},-\lambda_{2}^{2n}\psi_{2},\cdots,-\lambda_{n}^{2n}\psi_{n},-\lambda_{1}^{*2n}\phi_{1}^{*},\cdots,-\lambda^{*2n}\phi_{n}^{*})^{\mathrm{T}},$
where * denotes the complex conjugate and
$\psi_{j}=[c_{1}(c\lambda_{j}\mathrm{e}^{B_{j}}+(H_{j}-D_{j})\mathrm{e}^{-B_{j}})+c_{2}(-c\lambda_{j}\mathrm{e}^{-B_{j}}+(H_{j}+D_{j})\mathrm{e}^{B_{j}})]\mathrm{e}^{A},$
$\phi_{j}=[c_{1}((H_{j}-D_{j})\mathrm{e}^{B_{j}}+c\lambda_{j}\mathrm{e}^{-B_{j}})+c_{2}((H_{j}+D_{j})\mathrm{e}^{-B_{j}}-c\lambda_{j}\mathrm{e}^{B_{j}})]\mathrm{e}^{-A},$
$A=\mathrm{i}(ax+bt),~{}~{}D_{j}=2\mathrm{i}\delta c^{2}+\lambda_{j}^{2}\mathrm{i}-a\mathrm{i},$
$H_{j}=\sqrt{-4c^{4}\delta^{2}-4c^{2}\delta\lambda_{j}^{2}+4ac^{2}\delta-c^{2}\lambda_{j}^{2}-\lambda_{j}^{4}+2a\lambda_{j}^{2}-a^{2}},$
$B=\frac{H_{j}x}{2}+H_{j}t\Bigg{\{}\left(4\delta^{2}+3\delta+\frac{3}{8}\right)c^{4}+a^{2}+\left(\lambda_{j}^{2}+2\delta\right)a+\left(\lambda_{j}^{2}+\delta\right)^{2}$
$\quad+\left[\left(-4\delta-\frac{3}{2}\right)a-4\left(\frac{\lambda_{j}^{2}}{2}+\delta\right)\left(\delta+\frac{1}{4}\right)\right]c^{2}\Bigg{\}},$
where $c_{1},c_{2}$ are the real coefficients.
Subsequently, we firstly consider the first-order solution, namely $n=1$. Assume $H_{1}=H_{1,R}+\mathrm{i}H_{1,I}$ and $\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$, we can find that the first-order solution comprises the hyperbolic function parts $\cosh(V_{1,H}t+H_{1,R}x),~{}\sinh(V_{1,H}t+H_{1,R}x)$ and the trigonometric function parts $\cos(V_{1,T}t+H_{1,I}x),~{}\sin(V_{1,T}t+H_{1,I}x)$, where
$V_{1,H}$ $=W_{1}H_{1,R}-W_{2}H_{1,I},~{}V_{T}=W_{1}H_{1,I}+W_{2}H_{1,R},$
$W_{1}$ $=\frac{1}{4}\Bigg{[}32\left(c^{2}-\frac{1}{2}\right)^{2}\delta^{2}+\Bigl{(}24c^{4}-16\alpha_{1}^{2}c^{2}+16\beta_{1}^{2}c^{2}-32ac^{2}-8c^{2}-16\beta_{1}^{2}$
$+16\alpha_{1}^{2}+16a\Bigr{)}\delta+3c^{4}+\bigl{(}-4\alpha_{1}^{2}+4\beta_{1}^{2}-12a\bigr{)}c^{2}+8\alpha_{1}^{4}+\bigl{(}-48\beta_{1}^{2}+8a\bigr{)}\alpha_{1}^{2}$
$+8\beta_{1}^{4}-8a\beta_{1}^{2}+8a^{2}\Bigg{]},$
$W_{2}$ $=4\alpha_{1}\beta_{1}\Big{(}-2c^{2}\delta+2\delta-\frac{c^{2}}{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a\Big{)}.$
where the symbols $R$ and $I$ denote the real part and the imaginary part.
The above two function parts respectively describe the localization and periodicity of the wave’s transverse distribution. In other words, the first-order solution is composed of the soliton parts and the periodic parts. It’s obvious that the velocities of the two parts are $\frac{V_{1,H}}{H_{1,R}}$ and $\frac{V_{1,T}}{H_{1,I}}$, respectively. We can derive the first-order breather solution from Eq.(3) when $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$. When $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$, the first-order solution will be converted into a continuous soliton. The breather-to-soliton state transition will be discussed.
2.1 $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$
In this case, the first-order solution become the localized wave solution with breathing behavior. Under the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$, when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0, the general breather solution is obtained as shown in Fig.(1(a)). The Akhmediev breather solution which is localized in time and periodic in space is presented in Fig.(1(b)) when the condition $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0 holds.
Figure 1: The first-order breather solution. (a) The general breather solution with $\lambda_{1}=\frac{1}{3}+\frac{1}{2}\mathrm{i},a=\frac{3}{10},c=1,\delta=\frac{1}{10},c_{1}=c_{2}=2,b=-\frac{11}{50}$; (b) The Akhmediev breather solution with $\lambda_{1}=\frac{4}{5}+\frac{\sqrt{14}}{10}\mathrm{i},a=\frac{3}{5},c=1,\delta=-\frac{1}{5},c_{1}=c_{2}=1,b=-\frac{751}{1000}$.
2.2 $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$
In this case, the breather solution will be converted to the continuous soliton solution. We will discuss different types of the nonlinear converted waves.
The condition $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ can be expanded in the following form,
$-2c^{2}\delta+2\delta-\frac{1}{2}c^{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a=0.$ (4)
Given that the condition (4) is satisfied, we can obtain the multi-peak soliton as is shown in Fig.2 by choosing appropriate parameters in Eq.(3) when $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}\neq 0$. We can find that Fig.(2(a))-Fig.(2(d)) are the oscillation W-shaped soliton, while Fig.(2(e))-Fig.(2(h)) are the oscillation M-shaped soliton. The transformation between the two kinds of solitons can be finished by changing the sign of the imaginary part $\beta_{1}$ of the eigenvalue. Futhermore, it’s observed that when the real part $\alpha_{1}>0$ of the eigenvalue decreases, the number of wave peaks for the oscillation W-shaped and M-shaped soliton will increases.
Figure 2: The multi-peak soliton solution with $a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$. (a) The oscillation W-shaped soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) The oscillation W-shaped soliton with $\lambda_{1}=\frac{3}{10}-\frac{\sqrt{1}}{5}\mathrm{i}$; (d) The cross-sectional view of (c) at $t=0$; (e) The oscillation M-shaped soliton with $\lambda_{1}=\frac{2}{5}+\frac{\sqrt{11}}{10}\mathrm{i}$; (f) The cross-sectional view of (e) at $t=0$; (g) The oscillation M-shaped soliton with $\lambda_{1}=\frac{3}{10}+\frac{\sqrt{1}}{5}\mathrm{i}$; (h) The cross-sectional view of (g) at $t=0$.
Based on the condition (4), unlike the multi-peak soliton, we make another condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ holds. In this case, the periodic wave solution and anti-dark soliton solution are obtained. We can derive two different types of periodic wave solutions as is shown in Fig.3 by choosing appropriate parameters when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0. It is worth noting that the period of the periodic wave is $\frac{2\pi}{H_{I}}$, hence when the period extend to infinity, namely $H_{I}\to 0$, the periodic wave solution will be converted to the W-shape soliton solution as is presented in Fig.(4(a))-(4(b)). Subsequently, when $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0, the first-order solution can be degenerate into the anti-dark soliton solution which is a bell-shaped localized wave on the plane wave background in Fig.(4(c))-(4(d)).
Figure 3: The periodic wave solution with $a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$. (a) $\lambda=\frac{1}{2}+\frac{\sqrt{15}}{10}\mathrm{i}$; (b) The cross-sectional view of (a) at $t=0$; (c) $\lambda=\frac{7}{20}+\frac{3}{20}\mathrm{i}$; (d) The cross-sectional view of (d) at $t=0$.
Figure 4: (a)The W-shaped soliton degenerated from the periodic wave Fig.(3(a)) with $\lambda_{1}=\frac{\sqrt{35}}{10}-\frac{1}{2}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{50},c_{1}=c_{2}=1,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$; (c) The anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (d) The cross-sectional view of (c) at $t=0$.
3 Nonlinear wave interactions
In this section, the breather-to-soliton state transition of the second-order solution is discussed. In other words, we will explore the nonlinear converted wave interactions. We choose $n=2$, $H_{2}=H_{2,R}+\mathrm{i}H_{2,I}$ and $\lambda_{2}=\alpha_{2}+\mathrm{i}\beta_{2}$, then the second-order solution contains an additional hyperbolic function part $\cosh(V_{2,H}t+H_{2,R}x),~{}\sinh(V_{2,H}t+H_{2,R}x)$ and a trigonometric function part $\cos(V_{2,T}t+H_{2,I}x),~{}\sin(V_{2,T}t+H_{2,I}x)$ compared to the first-order solution. Hence, we need to consider two state transition conditions, namely $\frac{V_{j,H}}{H_{j,R}}=\frac{V_{j,T}}{H_{j,I}},j=1,2$ or $\frac{V_{j,H}}{H_{j,R}}\neq\frac{V_{j,T}}{H_{j,I}},j=1,2$.
Firstly, under the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}\neq\frac{V_{2,T}}{H_{2,I}}$, the interactions between the breather and other nonlinear converted waves can be obtained. Fig.5 exhibits the interaction between the breather and the anti-dark soliton when additional conditions $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ and $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)>$ 0 are satisfied. When $04\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<$ 0 holds, we can obtain the interaction between the breather and the periodic wave as is shown in Fig.6. Moreover, when the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ is satisfied, the second-order solution can be degenerated into the interaction between the breather and multi-peak soliton and we can see Fig.7.
Figure 5: The interaction between the breather and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{59}}{10}\mathrm{i},\lambda_{2}=1+\frac{1}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
Figure 6: The interaction between the breather and the periodic wave with $\lambda_{1}=\frac{7}{20}-\frac{3}{20}\mathrm{i},\lambda_{2}=1-\frac{1}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
Figure 7: The interaction between the breather and the multi-peak soliton with $\lambda_{1}=\frac{2}{5}-\frac{\sqrt{11}}{10}\mathrm{i},\lambda_{2}=-\frac{1}{2}+\mathrm{i},a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$; (b) The density plot for (a).
Subsequently, based on the conditions $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$ and $\frac{V_{2,H}}{H_{2,R}}=\frac{V_{2,T}}{H_{2,I}}$, when the conditions $\alpha_{j}^{2}-\beta_{j}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0,j=1,2$ hold and $24\beta_{j}^{2}\alpha_{j}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right),j=1,$ 2 show the same signs, we can obtain the interaction about the W-shaped soliton and the anti-dark soliton by adjusting the parameters. In Fig.8, the interaction between the W-shaped soliton and the W-shaped soliton is obtained. The interaction between the W-shaped soliton and the anti-dark soliton is presented in Fig.9. Meanwhile, we can also derive the interaction between the anti-dark soliton and the anti-dark soliton in Fig.LABEL:ffg3.
Figure 8: The interaction between the W-shaped soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}-\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
Figure 9: The interaction between the anti-dark soliton and the W-shaped soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=-\frac{4}{5}+\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
Figure 10: The interaction between the anti-dark soliton and the anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},\lambda_{2}=\frac{4}{5}-\frac{3\sqrt{6}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The cross-sectional view of (a) at $t=0$.
4 Double-pole breather-to-soliton transitions
In this section, we start with the double-pole solution of Eq.(1) given in [27] to investigate the double-pole breather-to-soliton transitions. It’s known that the double-pole solution can be rewritten as
$q[2]=q[0]-\dfrac{2\mathrm{i}|\Omega_{3}|}{|\Omega_{4}|},$ (5)
with
$\Omega_{4}$ $=(S_{1},\frac{d}{d\epsilon}S_{1},S_{2},\dfrac{d}{d\epsilon}S_{2})^{\mathrm{T}},~{}\Omega_{3}=(S_{3},\frac{d}{d\epsilon}S_{3},S_{4},\dfrac{d}{d\epsilon}S_{4})^{\mathrm{T}},$
$S_{1}$ $=\lim_{\epsilon\to 0}[(\lambda_{1}^{2}\psi_{1},\psi_{1},-\lambda_{1}^{4}\psi_{1},\lambda_{1}\phi_{1})^{\mathrm{T}}|_{\lambda_{1}=\lambda_{1}+\epsilon}],$
$S_{2}$ $=\lim_{\epsilon\to 0}[(\lambda_{1}^{*2}\phi_{1,*},\phi_{1}^{*},-\lambda_{1}^{*4}\phi_{1}^{*},-\lambda_{1}^{*}\psi_{1}^{*})^{\mathrm{T}}|_{\lambda_{1}=\lambda_{1}+\epsilon}],$
$S_{3}$ $=\lim_{\epsilon\to 0}[(\lambda_{1}^{2}\psi_{1},\psi_{1},-\lambda_{1}^{3}\phi_{1},\lambda_{1}\phi_{1})^{\mathrm{T}}|_{\lambda_{1}=\lambda_{1}+\epsilon}],$
$S_{4}$ $=\lim_{\epsilon\to 0}[(\lambda_{1}^{*2}\phi_{1,*},\phi_{1}^{*},-\lambda_{1}^{*3}\psi_{1}^{*},-\lambda_{1}^{*}\psi_{1}^{*})^{\mathrm{T}}|_{\lambda_{1}=\lambda_{1}+\epsilon}],$
where $\phi_{1},\psi_{1}$ are the same as Eq.(3).
Similarly, the state transition condition of the double-pole solution is $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$. Therefore, based on the condition, we will obtain the double-pole anti-dark soliton, double-pole multi-peak soliton and double-pole periodic wave solution. When the additional condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}\neq 0$ is satisfied, the double-pole solution can be reduced into the double-pole multi-peak soliton solution. We can see that Fig.11 presents the interaction between two series of the double-pole multi-peak soliton. Their wave peaks achieve the maximal amplitude at $(0,0)$. If the condition $\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)c^{2}-2a}{2}=0$ holds, when $4\beta_{1}^{2}\alpha_{1}^{2}+\frac{c^{4}}{4}+c^{2}\left(2c^{2}\delta-a\right)<0(>0)$, we can obtain the double-pole periodic wave solution and the double-pole anti-dark soliton solution as is shown in Fig.12 and Fig.13, respectively. The double-pole periodic wave solution possesses a periodic chained high-amplitude region which exhibits the double-pole phenomena.
Figure 11: (a) The double-pole multi-peak soliton solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{5}}{5}\mathrm{i},a=\frac{2}{5},c=1,\delta=\frac{7}{10},c_{1}=c_{2}=2,b=-\frac{5109}{1000}$; (b) The density plot for (a).
Figure 12: (a) The double-pole periodic wave solution with $\lambda_{1}=\frac{1}{2}-\frac{\sqrt{15}}{5}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
Figure 13: (a) The double-pole anti-dark soliton with $\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i},a=\frac{3}{10},c=1,\delta=-\frac{3}{20},c_{1}=c_{2}=2,b=-\frac{1}{160}$; (b) The density plot for (a).
We will apply the asymptotic analysis method described in [28] to explore the asymptotic behavior of the double-pole anti-dark soliton. For convenience, assume $c=1,c_{1}=c_{2}=2,\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$ that under the corresponding state transition conditions, we can obtain $a=\alpha_{1}^{2}-\beta_{1}^{2}+\frac{4\delta+1}{2},\delta=-\frac{3\alpha_{1}^{2}}{2}+\frac{3\beta_{1}^{2}}{2}$ and $14\beta_{1}^{2}>$ 1. Hence, the double-pole anti-dark soliton solution is
$q_{d}=\dfrac{3072e^{\mathrm{i}D_{1}}(E_{1}e^{\frac{D_{2}\Delta}{8}}+E_{2}e^{\frac{3D_{2}\Delta}{8}}+E_{3}e^{\frac{D_{2}\Delta}{2}}-\frac{1}{2}E_{4}e^{\frac{D_{2}\Delta}{4}}-E_{5})}{G_{1}e^{\frac{D_{2}\Delta}{2}}+G_{2}e^{\frac{D_{2}\Delta}{4}}+384\beta(G_{3}e^{\frac{D_{2}\Delta}{8}}+G_{4}e^{\frac{3D_{2}\Delta}{8}}-G_{5})},$ (6)
where
$D_{1}$ $=\frac{25(\beta_{1}-\alpha_{1})^{3}(\beta_{1}+\alpha_{1})^{3}t}{4}+2\left(\beta_{1}^{2}-\alpha_{1}^{2}+\frac{1}{4}\right)x,$
$D_{2}$ $=\left(6\beta_{1}^{4}+4(5\alpha_{1}^{2}-1)\beta_{1}^{2}+6\alpha_{1}^{4}+4\alpha_{1}^{2}+1\right)t-4x,~{}\Delta=\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)},$
and $E_{j},j=1,\cdots,5$ and $G_{j},j=1,\cdots,5$ are too long to be included here, so we omit them.
Subsequently, we substitute $x=\frac{\left(6\beta_{1}^{4}+4(5\alpha_{1}^{2}-1)\beta_{1}^{2}+6\alpha_{1}^{4}+4\alpha_{1}^{2}+1\right)t-D_{2}}{4}$ into Eq.(6) so that the double-pole solution is only related to $D_{2}$ and $t$. Numerous articles about the asymptotic analysis method have illustrated that the characteristic line of the double-pole solution is curved[29, 30, 31]. Therefore, according to the relation $e^{D_{2}}\geq D_{2}(D_{2}\to+\infty)$ and $e^{D_{2}}\leq D_{2}(D_{2}\to-\infty)$, there must exist an asymptotic balance relation between $e^{D_{2}}$ and $t$. The explicit expressions of the asymptotic solitons are exhibited as follows.
$\bullet$  When $e^{D_{2}}\sim t$, $D_{2}\to+\infty$,
$[u]_{\mathrm{I}}^{\pm}\to e^{D_{1}}\Bigg{[}1+\frac{K_{1}}{2K_{2}\cosh\Big{(}\frac{D_{2}\Delta}{8}+\ln K_{3}\Big{)}+384\beta_{1}K_{4}}\Bigg{]},~{}~{}t\to\pm\infty,$ (7)
where
$K_{1}$ $=32t\cdot(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3},$
$\quad\cdot\left(4i\beta_{1}^{2}-i-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right),$
$K_{2}$ $=32\sqrt{2}\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}$
$\quad\cdot\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}},$
$K_{3}$ $=\frac{\beta_{1}^{4}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)(\alpha_{1}^{2}+\beta_{1}^{2})\sqrt{2}}{2\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}}$
$\quad\cdot\frac{1}{\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}}},$
$K_{4}$ $=\frac{t(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\cdot\left(4i\beta_{1}^{2}-i-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right).$
$\bullet$  When $e^{D_{2}}\sim t$, $D_{2}\to-\infty$,
$[u]_{\mathrm{II}}^{\pm}\to e^{D_{1}}\Bigg{[}1+\frac{L_{1}}{2L_{2}\cosh\Big{(}\frac{D_{2}\Delta}{8}+\ln L_{3}\Big{)}+384\beta_{1}L_{4}}\Bigg{]},~{}~{}t\to\pm\infty,$ (8)
where
$L_{1}$ $=32t(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3}$
$\quad\cdot\left(4i\beta_{1}^{2}-i+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right),$
$L_{2}$ $=32\sqrt{2}\sqrt{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}$
$\quad\cdot\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)},$
$L_{3}$ $=-\frac{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{2}\beta_{1}^{4}\sqrt{2}}{\sqrt{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}},$
$\quad\cdot\frac{1}{\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)}},$
$L_{4}$ $=\frac{t(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\cdot\left(4i\beta_{1}^{2}-i+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right).$
We can find that the characteristic lines of the asymptotic solitons above are $\frac{D_{2}\Delta}{8}+\ln K_{3}=0$ and $\frac{D_{2}\Delta}{8}+\ln L_{3}=0$, both contain the logarithmic terms which can reflect the double-pole properties. When $a=\frac{3}{10},c=1,\delta=-\frac{3}{20},\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i}$ are taken, we can find that the characteristic trajectories of the corresponding asymptotic solitons in Fig.14 fit perfectly with the contour plot of Fig.13. Furthermore, we compare the asymptotic solitons (7), (8) and the exact solution (5) when the parameters are given. We can find that the asymptotic soliton perfectly matches the exact solution in the far-field region as is shown in Fig.15. This demonstrates the validity of our asymptotic analysis.
Figure 14: The contour plot for the double-pole anti-dark soliton.
(a) $t=30$
(b) $t=-30$
(c) $t=60$
(d) $t=-60$
Figure 15: (a)The comparison of the asymptotic soliton and the exact solution (black line) for the double-pole anti-dark soliton. (a)(c) Red: the asymptotic soliton $[u]_{\mathrm{I}}^{+}$; Green: the asymptotic soliton $[u]_{\mathrm{II}}^{+}$; (b)(d) Red: the asymptotic soliton $[u]_{\mathrm{I}}^{-}$; Green: the asymptotic soliton $[u]_{\mathrm{II}}^{-}$.
5 Conclusions and discussions
In conclusion, we explore the breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation.
$\bullet$  We obtain different types of the nonlinear converted waves, such as the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution.
$\bullet$  The interactions between various noninear converted waves are exhibited.
$\bullet$  The double-pole breather-to-soliton transitions are studied and the asymptotic analysis method is applied to analyze the asymptotic behavior of the double-pole anti-dark soliton.
Our work will provide a reference for future projects about the breather-to-soliton transitions.
6 Data availability
No data was used for the research described in the article.